# GTPBD: A Fine-Grained Global Terraced Parcel and Boundary Dataset

### v1

@@ Abstract
Agricultural parcels serve as basic units for conducting agricultural practices and applications, which is vital for land ownership registration, food security assessment, soil erosion monitoring, etc. However, existing agriculture parcel extraction studies only focus on mid-resolution mapping or regular plain farmlands while lacking representation of complex terraced terrains due to the demands of precision agriculture. In this paper, we introduce a more fine-grained terraced parcel dataset named GTPBD (Global Terraced Parcel and Boundary Dataset), which is the first fine-grained dataset covering major worldwide terraced regions with more than 200,000 complex terraced parcels with manually annotation. GTPBD comprises 47,537 high-resolution images with three-level labels, including pixel-level boundary labels, mask labels, and parcel labels. It covers seven major geographic zones in China and transcontinental climatic regions around the world. Compared to the existing datasets, the GTPBD dataset brings considerable challenges due to the: (1) terrain diversity; (2) complex and irregular parcel objects; and (3) multiple domain styles. Our proposed GTPBD dataset is suitable for four different tasks, including semantic segmentation, edge detection, terraced parcel extraction and unsupervised domain adaptation (UDA) tasks. Accordingly, we benchmark the GTPBD dataset on eight semantic segmentation methods, four edge extraction methods, three parcel extraction methods and five UDA methods, along with a multi-dimensional evaluation framework integrating pixel-level and object-level metrics. GTPBD fills a critical gap in terraced remote sensing research, providing a basic infrastructure for fine-grained agricultural terrain analysis and cross-scenario knowledge transfer. The code and data are available at https://github.com/Z-ZW-WXQ/GTPBG/.
@@ 1 Introduction
In ancient China, terraced fields were given the reputation of "Dragon Spines" ("{CJK}UTF8gbsn龙脊" in Chinese), which is not only a poetic description of their winding and magnificent forms, but also a high praise in agricultural civilization. Nowadays, approximately 120 million acres of terraced fields worldwide still support more than 500 million mountainous populations unicef2024state , reducing soil erosion by more than 23.7 billion tons annually, which indicates significant ecological and economic values ferro2017terrace . Therefore, a better and fine-grained understanding of terraced area is needed for a sustainable agricultural ecosystem, which is vital for land ownership registration, food security assessment, etc spencer1961origin ; arnaez2015effects .
With the development of artificial intelligence techniques, terrace mapping have progressed significantly, from pixel-based and object-oriented methods to machine learning and deep learning-based methods from remote sensing images. Existing terrace mapping only focus on mid-resolution mapping cao202030 ; li202510 ; ding2021evaluation using Sentinel-1/2, Landsat, MODIS, etc. Meanwhile, some researchers have collected public large-scale agricultural parcels and boundaries datasets from high-resolution remote sensing images. To improve the agricultural parcel extraction, advanced deep learning methods have been used in parcel-level extraction and vectorization yu2022research ; pan2023e2evap ; ReauNet-Sober ; SEANet . Despite the impressive capabilities demonstrated by deep learning for agricultural parcel extraction, fine-grained terraced parcel and boundary benchmarks for agricultural remote sensing remain scarce due to they are extremely complex and unexpectedly irregular. Existing agricultural parcel dataset primarily focus on scenarios of regular plain farmlands (See Fig. 1 (a)). Datasets designed for agricultural parcel dataset from high-resolution remote sensing images exhibit notable limitations, primarily in terms of terrain diversity, annotation difficulty and multiple tasks, as shown in Fig. 1.
@@
In response to the aforementioned issues, this paper proposes a more challenging terraced parcel dataset named GTPBD (Global Terraced Parcel and Boundary Dataset), which is the first fine-grained dataset covering major worldwide terraced regions with more than 200,000 complex terraced parcels with manually annotation. In summary, the contributions of this work can be concluded as:
@@
Table 1: Comparison between GTPBD and the main agricultural field datasets, including raster mask datasets and vector-only datasets.
Dataset Source Images Area (km2) Resolution Classes Global Task (m) coverage SS APE ED UDA Raster mask FHAPD FHAPD/HBGNet GaoFen-1/2 68,982 $<$500 0.5 2 FTW FTW Sentinel-2 70,462 166,293 10 2 AI4B (S-2) AI4Boundaries Sentinel-2 7,831 51,321 10 2 AI4B (Ortho) AI4Boundaries Aerial 7,598 1,992 1 2 PASTIS PASTIS Sentinel-2 2,433 3,986 10 19 CP-Set cp-set GF-2 949 249 1 2 GFSAD30 GFSAD30 Landsat-8 64,800 18,740,000 30 2 Vector-only GloCAB GloCAB Sentinel-2 190,832 N/A 10 2 EuroCropsML Eurocropsml Sentinel-2 706,683 >4,000 10 176 AI4SmallFarms AI4SmallFarms Sentinel-2 62 1,550 10 2 India Smallholder Airbus SPOT-6/7 10,000 30-50 1.5-4.8 2 Boundaries India & PlanetScope GTPBD (Ours) GF-2 & Google Earth 47,537 885 0.1–1 3
@@ 2.1 Agricultural parcel and boundary datasets
In recent years, significant efforts have been made to establish comprehensive agricultural parcel datasets (see Table 1). Early datasets such as GFSAD30 GFSAD30 offer extensive global coverage but at a relatively coarse resolution of 30 meters, limiting their suitability for detailed parcel delineation tasks. PASTIS PASTIS and FTW FTW developed from Sentinel-2 imagery with resolutions around 10 meters, advanced the granularity of available agricultural data. FHAPD FHAPD/HBGNet and CP-Set cp-set comprise high-resolution GaoFen-2 (GF-2) imagery at 0.5 and 1 meters resolution, enhancing the spatial detail for agricultural parcel extraction tasks. But they predominantly include regularized, planar agricultural fields, limiting their applicability to terraced landscapes. AI4Boundaries AI4Boundaries provides annotations suitable for tasks such as semantic segmentation and edge detection, yet these efforts still primarily target flat agricultural areas, with limited consideration of terraced agricultural systems. Besides, vector-only datasets such as GloCAB GloCAB and EuroCropsML Eurocropsml though comprising a large number of images, still fall short of achieving true global coverage, and their utility is restricted to agricultural parcel extraction (APE) tasks because they only provide vector annotations. India Smallholder Boundaries India despite offering relatively high spatial resolution, focuses on flat farmlands in India and supports only APE tasks, making it less suitable for broader situation.
@@ 2.2 Remote sensing tasks
Agricultural Parcel Extraction (APE) delineates farmland units for precision agriculture APEusage1 ; APEusage2 . Most CNN-based methods treated APE as single-task segmentation, using encoder–decoder backbones to predict parcel masks but often yielding unclosed or jagged boundaries in heterogeneous fields APEdisadvantage . To address this, edge-aware multi-task architectures have emerged: SEANet SEANet jointly predicts masks, distances, and edges to enforce closed parcels, REAUNet ReauNet-Sober integrates Sobel filters and attention for multi-scale edge enhancement, HBGNet FHAPD/HBGNet extracts low-level boundary features and high-level parcel semantics in parallel with a dual-branch framework. Despite these models perform well on planar benchmarks like FHAPD FHAPD/HBGNet and AI4Boundaries AI4Boundaries , their robustness to terraced ridge geometries is untested. We therefore benchmark them on GTPBD to provide systematic evaluation of APE methods in complex terraced agricultural landscapes.
@@
Unsupervised Domain Adaptation (UDA) aims to transfer knowledge from a labeled source domain to an unlabeled target domain UDA1 . UDA methods for remote sensing can be broadly grouped into adversarial training and self-training approaches. Adversarial training such as Cycada cycada and CLAN CLAN leverage discriminators to align feature distributions, they incur complex optimization overhead and may not directly address style discrepancies. Representative self-training frameworks include DAFormer DAFormer , which integrates rare-class sampling and pseudo-label denoising. In addition, input-level adaptation methods like FDA FDA perform only low-frequency transfer in the Fourier domain, requiring no adversarial network or pseudo label. Although these UDA methods have been widely used in remote sensing community with some improvements, existing public remote sensing datasets have been limited by urban scenes li2024hyunida ; chen2025ban or land cover mapping wang2loveda ; zheng2024open . To this end, the GTPBD dataset is proposed for a more challenging benchmark, promoting future research of cross-regional or temporal agricultural and terraced parcel extraction algorithms and its applications.
@@ 3.1 Image collection and manual annotation
Spatial Distribution We give selection priority to regions with dense terrace distribution and high topographic complexity according to GTM li202510 . We also select typical terraces in diverse landforms, such as the alpine terraces in southwestern China and the Sa Pa terraces in Vietnam, to evaluate the model’s generalization ability across varied landforms. Fig 2 displays the major study areas covering fourteen countries around the world, including China, Vietnam, Tunisia, Ethiopia, Peru, Mexico and more. In addition, our dataset covers seven major geographic regions of China, including Central China, South China, North China, East China, Northwest China, Northeast China, and Southwest China. More images could be found in Appendix B.1.
Image Resolution We download a series of high-resolution terraced field images from Google Earth, with 0.1$\sim$1m, ensuring the delicacy and clarity of the images. More importantly, these high-resolution terraced field images not only cover diverse seasonal changes but also capture the scenery under different lighting conditions, thus comprehensively showcasing the appearance of terraced fields in different environments. It is worth mentioning that all of these images were carefully selected from cloud-free terraced fields between 2021 and 2025, providing extremely valuable and high-quality data resources for related research.
@@
Three-Level Labels As for mask labels, we use GDAL’s rasterize function for fully connected rasterization (all touched strategy), establish a binary semantic segmentation benchmark, and assign 1 to terraced areas and 0 to backgrounds. As for boundary labels, we use a 3$\times$3 rectangular grid mask to perform a single morphological etching and construct a 3-pixel-wide edge detection dedicated label. As for parcel labels, we generate pure parcel regions through mask boundary using the XOR operation to achieve terraced parcel extraction tasks. Appendix B.2 displays the differences between previous dataset and our GTPBD.
Dataset Construction We crop the original remote sensing image with the size of 512$\times$512 pixels according to existing works AI4Boundaries . We divide the GTPBD dataset into training set, validation set, and testing set by the raito of 60%, 20% and 20% (see Fig. 3 (b)), which are used for model training, tuning and hyperparameter selection, and performance evaluation, respectively.
@@ 3.3 Differences among three domains
We seperate our GTPBD into three domains: South China (South), North China (North) and other regions except China (Global). The detailed statics of these three domains are shown in Fig. 4 (a) and (b). According to Fig. 5, the characteristics from different regions are really different. For the spectral statistics, the mean values are similar (Fig. 4 (c)) because of the large-scale homogeneous geographical areas and diverse land cover types. We could observe that the South domain have lower standard deviations. As is shown in Fig. 4 (d), although all domains exist “long tail” phenomenon (logarithmic scale in vertical axis), the South domain has the most of the parcels have relatively small scales. When faced with large-scale terraced parcel and boundary extraction tasks, the differences between different scenes or regions bring new challenges to the model generalization and transferability.
@@ 4.1 Experimental setup
The data splits followed Sec. 3.1. During the training, we used the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay of $10^{-4}$. For the data augmentation, 512$\times$512 patches were randomly cropped with random mirroring and rotation. We implemented all our experiments on NVIDIA RTX 4090 GPU. We benchmark the GTPBD dataset on eight semantic segmentation (U-Net unet , DeepLabV3 deeplabv3 , PSPNet pspnet , NonLocal nonlocal , OCRNet OCRNet , K-Net knet , SegFormer segformer and Mask2Former mask2former ) methods, four edge extraction methods (MuGE Muge , PiDiNet pidinet , UEAD UEAD and REAUNet-Sober ReauNet-Sober ), three parcel extraction methods (REAUNet ReauNet-Sober , SEANet SEANet , HBGNet FHAPD/HBGNet ) and five UDA methods (Source only, FDA FDA , PiPa PiPa , HRDA HRDA and DAFormer DAFormer ). More implementation details are provided in the Appendix C.
@@ 4.2 Evaluation metrics
Table 2: Comparisons of different semantic segmentation methods on the GTPBD dataset (%). More visualization results and the performance of regions could be found in Appendix E.1.2 and E.1.1.
@@
U-Net unet 78.21 72.79 60.50 74.18 75.40
Deeplabv3 deeplabv3 81.57 81.64 68.93 79.99 81.61
PSPNet pspnet 79.03 84.77 69.20 79.48 81.80
Nonlocal nonlocal 82.67 80.22 68.67 80.10 81.43
OCRNet OCRNet 77.53 86.22 68.98 78.92 81.65
K-Net knet 78.69 81.18 66.55 77.81 79.92
SegFormer segformer 80.48 83.03 69.11 79.82 81.73
Mask2Former mask2former 73.02 94.25 69.91 77.94 82.29
@@ 4.3 Semantic Segmentation Results
Semantic segmentation plays a vital role in agricultural parcel extraction. As summarized in Table 3, these models exhibit notable discrepancies in segmentation performance, reflecting varying capabilities in modeling fine-grained agricultural boundaries. Specifically, the NonLocal model achieved the highest Prec. (82.67%) and OA (80.10%), indicating strong discrimination with minimal false positives. In contrast, Mask2Former demonstrated superior generalization, attaining the best scores in Rec. (94.25%), IoU (69.91%), and F1-score (82.29%). These findings highlight a trade-off between precision and recall among different architectures and emphasize the robustness of transformer-based models under complex terraced scenarios. Fig. 6 visually compare the segmentation results on a sample patch, with red highlighting false positives and blue highlighting false negatives. More visualization results and the performance of regions could be found in Appendix E.1.2 and E.1.1.

### v3

@@ Abstract
Agricultural parcels serve as basic units for conducting agricultural practices and applications, which is vital for land ownership registration, food security assessment, soil erosion monitoring, etc. However, existing agriculture parcel extraction studies only focus on mid-resolution mapping or regular plain farmlands while lacking representation of complex terraced terrains due to the demands of precision agriculture. In this paper, we introduce a more fine-grained terraced parcel dataset named GTPBD (Global Terraced Parcel and Boundary Dataset), which is the first fine-grained dataset covering major worldwide terraced regions with more than 200,000 complex terraced parcels with manually annotation. GTPBD comprises 47,537 high-resolution images with three-level labels, including pixel-level boundary labels, mask labels, and parcel labels. It covers seven major geographic zones in China and transcontinental climatic regions around the world. Compared to the existing datasets, the GTPBD dataset brings considerable challenges due to the: (1) terrain diversity; (2) complex and irregular parcel objects; and (3) multiple domain styles. Our proposed GTPBD dataset is suitable for four different tasks, including semantic segmentation, edge detection, terraced parcel extraction and unsupervised domain adaptation (UDA) tasks. Accordingly, we benchmark the GTPBD dataset on eight semantic segmentation methods, four edge extraction methods, three parcel extraction methods and five UDA methods, along with a multi-dimensional evaluation framework integrating pixel-level and object-level metrics. GTPBD fills a critical gap in terraced remote sensing research, providing a basic infrastructure for fine-grained agricultural terrain analysis and cross-scenario knowledge transfer. The code and data are available at https://github.com/Z-ZW-WXQ/GTPBD/.
@@ 1 Introduction
In ancient China, terraced fields were given the reputation of ”Dragon Spines” (”龙脊” in Chinese), which is not only a poetic description of their winding and magnificent forms, but also a high praise in agricultural civilization. Nowadays, approximately 120 million acres of terraced fields worldwide still support more than 500 million mountainous populations unicef2024state, reducing soil erosion by more than 23.7 billion tons annually, which indicates significant ecological and economic values ferro2017terrace. Therefore, a better and fine-grained understanding of terraced area is needed for a sustainable agricultural ecosystem, which is vital for land ownership registration, food security assessment, etc spencer1961origin; arnaez2015effects.
With the development of artificial intelligence techniques, terrace mapping have progressed significantly, from pixel-based and object-oriented methods to machine learning and deep learning-based methods from remote sensing images. Existing terrace mapping only focus on mid-resolution mapping cao202030; li202510; ding2021evaluation using Sentinel-1/2, Landsat, MODIS, etc. Meanwhile, some researchers have collected public large-scale agricultural parcels and boundaries datasets from high-resolution remote sensing images. To improve the agricultural parcel extraction, advanced deep learning methods have been used in parcel-level extraction and vectorization yu2022research; pan2023e2evap; ReauNet-Sober; SEANet.
Despite the impressive capabilities demonstrated by deep learning for agricultural parcel extraction, fine-grained terraced parcel and boundary benchmarks for agricultural remote sensing remain scarce due to they are extremely complex and unexpectedly irregular. Existing agricultural parcel dataset primarily focus on scenarios of regular plain farmlands (See Fig. 1 (a)). Datasets designed for agricultural parcel dataset from high-resolution remote sensing images exhibit notable limitations, primarily in terms of terrain diversity, annotation difficulty and multiple tasks, as shown in Fig. 1.
@@
In response to the aforementioned issues, this paper proposes a more challenging terraced parcel dataset named GTPBD (Global Terraced Parcel and Boundary Dataset), which is the first fine-grained dataset covering major worldwide terraced regions with more than 200,000 complex terraced parcels with manually annotation.
In summary, the contributions of this work can be concluded as:
@@
Table 1: Comparison between GTPBD and the main agricultural field datasets.
Dataset Source Images Area (km2) Resolution Classes Global Task (m) Coverage SS APE ED UDA FHAPD FHAPD/HBGNet GaoFen-1/2 68,982 <<1000 1-2 2 GTM li202510 Sentinel-2 108,300 853,161 10 2 FTW FTW Sentinel-2 70,462 166,293 10 2,3 AI4Boundaries (Sentinel-2) AI4Boundaries Sentinel-2 7,831 51,321 10 2 AI4Boundaries (Ortho) AI4Boundaries Aerial 7,598 1,992 1 2 PASTIS PASTIS Sentinel-2 2,433 3,986 10 19 CropHarvest tseng2021cropharvest Sentinel-1/2 95,186 ¿5,000,000 10-60 348 GFSAD30 GFSAD30 Landsat-8 64,800 18,740,000 30 2 GloCAB GloCAB Sentinel-2 190,832 N/A 10 2 AI4SmallFarms AI4SmallFarms Sentinel-2 62 1,550 10 2 India Smallholder Boundaries India Airbus SPOT-6/7 & PlanetScope 10,000 30-50 1.5-4.8 2 GTPBD (Ours) GF-2 & Google Earth 47,537 885 0.5–0.7 3
@@ 2.1 Agricultural parcel and boundary datasets
In recent years, considerable progress has been made in developing agricultural parcel and boundary datasets (see Table 1). Early datasets such as GFSAD30 GFSAD30 provide global coverage but at a coarse resolution of 30 m, making them unsuitable for fine-grained parcel delineation. More recent efforts like PASTIS PASTIS and FTW FTW, built on Sentinel-2 imagery (10 m resolution), improved spatial granularity but still mainly represent regular, planar agricultural fields. Higher-resolution datasets such as FHAPD FHAPD/HBGNet and AI4Boundaries(Ortho) AI4Boundaries offer imagery at 1 m resolution, enhancing spatial detail. However, these remain limited to structured flat farmland, with little focus on irregular terraced landscapes. Importantly, none of the existing datasets support unsupervised domain adaptation (UDA), despite its necessity for cross-region generalization.
@@ 2.2 Remote sensing tasks
Agricultural Parcel Extraction (APE) delineates farmland units for precision agriculture APEusage1; APEusage2. Most CNN‐based methods treated APE as single‐task segmentation, using encoder–decoder backbones to predict parcel masks but often yielding unclosed or jagged boundaries in heterogeneous fields APEdisadvantage. To address this, edge‐aware multi‐task architectures have emerged: SEANet SEANet jointly predicts masks, distances, and edges to enforce closed parcels, REAUNet ReauNet-Sober integrates Sobel filters and attention for multi‐scale edge enhancement, HBGNet FHAPD/HBGNet extracts low-level boundary features and high-level parcel semantics in parallel with a dual-branch framework. Despite these models perform well on planar benchmarks like FHAPD FHAPD/HBGNet and AI4Boundaries AI4Boundaries, their robustness to terraced ridge geometries is untested. We therefore benchmark them on GTPBD to provide systematic evaluation of APE methods in complex terraced agricultural landscapes.
@@
Unsupervised Domain Adaptation (UDA) aims to transfer knowledge from a labeled source domain to an unlabeled target domain UDA1. UDA methods for remote sensing can be broadly grouped into adversarial training and self-training approaches. Adversarial training such as Cycada cycada and CLAN CLAN leverage discriminators to align feature distributions, they incur complex optimization overhead and may not directly address style discrepancies. Representative self‐training frameworks include DAFormer DAFormer, which integrates rare‐class sampling and pseudo‐label denoising. In addition, input‐level adaptation methods like FDA FDA perform only low‐frequency transfer in the Fourier domain, requiring no adversarial network or pseudo label. Although these UDA methods have been widely used in remote sensing community with some improvements, existing public remote sensing datasets have been limited by urban scenes li2024hyunida; chen2025ban or land cover mapping wang2loveda; zheng2024open. To this end, the GTPBD dataset is proposed for a more challenging benchmark, promoting future research of cross-regional or temporal agricultural and terraced parcel extraction algorithms and its applications.
@@ 3.1 Image collection and manual annotation
Spatial Coverage and Resolution Following GTM li202510, we prioritized regions with dense terrace distribution and complex topography, and further included representative terraces across diverse landforms—such as alpine terraces in southwestern China and the Sa Pa terraces in Vietnam—to assess model generalization across varied terrains. The dataset spans fourteen countries worldwide, including China, Vietnam, Tunisia, Ethiopia, Peru, and Mexico, as well as seven major geographic regions of China (Central, South, North, East, Northwest, Northeast, and Southwest). High-resolution images were collected from GF-2 and Google Earth with spatial resolutions ranging from 0.1 m to 1 m, covering different seasons and lighting conditions. All images were selected from cloud-free scenes acquired between 2021 and 2025 to ensure consistency and research suitability. Examples are shown in Fig.2, with additional samples in AppendixB.1.
@@
Three-Level Labels As for mask labels, we use GDAL’s rasterize function for fully connected rasterization (all touched strategy), establish a binary semantic segmentation benchmark, and assign 1 to terraced areas and 0 to backgrounds. As for boundary labels, we use a $\times3$ rectangular grid mask to perform a single morphological etching and construct a 3-pixel-wide edge detection dedicated label. As for parcel labels, we generate pure parcel regions through mask boundary using the XOR operation to achieve terraced parcel extraction tasks. Appendix B.2 displays the differences between previous dataset and our GTPBD.
Dataset Construction To address potential spatial correlation or leakage between patches, we first resampled all source images and corresponding labels to a unified resolution of 0.5–0.7 m. Images from the same geographic region share the same target resolution and were split into training (60%), validation (20%), and testing (20%) sets before cropping, ensuring spatial independence between subsets. Following common practice in high-resolution agricultural benchmarks such as AI4BoundariesAI4Boundaries, Agriculture-Vision 9157397, and FUSU NEURIPS2024_eeb3df2d, we adopted a patch size of 512 × 512 pixels. Compared to 256 × 256 patches, this size preserves more complete geometric patterns and semantic context, such as parcel shapes and boundaries.
@@ 3.3 Differences among three domains
We seperate our GTPBD into three domains: South China (South), North China (North) and other regions except China (Global). The detailed statics of these three domains are shown in Fig. 4 (a) and (b). According to Fig. 5, the characteristics from different regions are really different.
For the spectral statistics, the mean values are similar (Fig. 4 (c)) because of the large-scale homogeneous geographical areas and diverse land cover types. We could observe that the South domain have lower standard deviations. As is shown in Fig. 4 (d), although all domains exist “long tail” phenomenon (logarithmic scale in vertical axis), the South domain has the most of the parcels have relatively small scales. When faced with large-scale terraced parcel and boundary extraction tasks, the differences between different scenes or regions bring new challenges to the model generalization and transferability.
@@ 4.1 Experimental setup
The data splits followed Sec. 3.1. During the training, we used the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay of $10^{-4}$. For the data augmentation, $\times512$ patches were randomly cropped with random mirroring and rotation. We implemented all our experiments on NVIDIA RTX 4090 GPU. We benchmark the GTPBD dataset on eight semantic segmentation methods (U-Net unet, DeepLabV3 deeplabv3, PSPNet pspnet, NonLocal nonlocal, OCRNet OCRNet, K-Net knet, SegFormer segformer and Mask2Former mask2former), four edge extraction methods (MuGE Muge, PiDiNet pidinet, UEAD UEAD and REAUNet-Sober ReauNet-Sober), three parcel extraction methods (REAUNet ReauNet-Sober, SEANet SEANet, HBGNet FHAPD/HBGNet) and five UDA methods (Source only, FDA FDA, PiPa PiPa, HRDA HRDA and DAFormer DAFormer). More implementation details are provided in the Appendix C.
@@ 4.2 Evaluation metrics
Table 2: Comparisons of different semantic segmentation methods on the GTPBD dataset (%). More visualization results and the performance of regions could be found in Appendix E.1.3 and E.1.2.
@@
UNet unet 74.11 54.93 46.09 75.46 63.09
DeepLabV3 deeplabv3 69.64 73.45 57.04 78.28 71.58
PSPNet pspnet 68.33 72.41 54.22 76.65 72.31
Nonlocal nonlocal 75.06 70.27 51.48 79.52 72.58
OCRNet OCRNet 73.05 71.56 53.87 76.95 72.30
K-Net knet 74.56 61.04 50.52 77.17 67.13
SegFormer segformer 74.45 69.07 55.84 78.14 71.66
Mask2Former mask2former 71.22 74.33 57.16 78.73 72.74
@@ 4.3 Semantic Segmentation Results
Semantic segmentation plays a vital role in agricultural parcel extraction. As summarized in Table 3, these models exhibit notable discrepancies in segmentation performance, reflecting varying capabilities in modeling fine-grained agricultural boundaries. Specifically, the NonLocal model achieved the highest Prec. (75.06%) and OA (79.52%), indicating strong discrimination with minimal false positives. In contrast, Mask2Former demonstrated superior generalization, attaining the best scores in Rec. (74.33%), IoU (57.16%), and F1-score (72.74%). These findings highlight a trade-off between precision and recall among different architectures and emphasize the robustness of transformer-based models under complex terraced scenarios. Fig. 6 visually compare the segmentation results on a sample patch, with red highlighting false positives and blue highlighting false negatives. As can be observed, conventional semantic segmentation can only distinguish farmland regions from non-farmland areas, but fails to delineate precise parcel boundaries — especially for adjacent parcels that share common borders. More visualization results and the performance of regions could be found in Appendix E.1.